OpenStudy (cp9454):

if a,b,c are three distinct real number in geometric progression, and a+b+c = xb , then x belongs in @camerondoherty

OpenStudy (cp9454):

ya nice try @eyedulla

OpenStudy (camerondoherty):

XD LMAO

OpenStudy (cp9454):

i didn't get the second one btw @eyedulla lol

OpenStudy (mathstudent55):

2, 6, 18 2, 8, 32 26 42 26 = x * 6 42 = x * 8 x = 26/6 = 13/3 x = 42/8 = 21/4 k, kd, kd^2 k + kd + kd^2 = k(1 + d + d^2) k(1 + d + d^2) = xkd x = (k(1 + d + d^2))/kd x = (1 + d + d^2)d

OpenStudy (mathstudent55):

Example 1 Example 2 2, 6, 18 2, 8, 32 Sum = 26 Sum = 42 26 = x * 6 42 = x * 8 x = 26/6 = 13/3 x = 42/8 = 21/4 In general k, kd, kd^2 Sum = k + kd + kd^2 = k(1 + d + d^2) k(1 + d + d^2) = xkd x = (k(1 + d + d^2))/kd x = (1 + d + d^2)/d

OpenStudy (mathstudent55):

What is the question?

OpenStudy (cp9454):

you have to find range in which x belongs

ganeshie8 (ganeshie8):

xd = 1 + d + d^2 d^2 + d(1-x) + 1 = 0 for d to be real, the discriminant needs to be non negative : (1-x)^2 - 4 >= 0 solve x

OpenStudy (cp9454):

Thats 100% right. thanks